OpenStudy (anonymous):

Which of the following expressions are equivalent? Justify your reasoning. 4√x3 1 x−1 10√x5•x4•x2 x 1 3 •x 1 3 •x 1 3

OpenStudy (decentnabeel):

which expression

OpenStudy (anonymous):

So the options are A. \[\sqrt[4]{x^3}\] B. \[\frac{ 1 }{ x^-1}\] C.\[\sqrt[10]{x^5*x^4*x^2}\] D.\[x^{1/3}\] * \[x^{1/3}\] * \[x^{1/3}\]

OpenStudy (anonymous):

@DecentNabeel hey can you help me out?

OpenStudy (decentnabeel):

ok these are option so what is the expression

OpenStudy (anonymous):

there is none, they're only asking which are equivalent

OpenStudy (anonymous):

so out of all the options, which ones equal the same thing

OpenStudy (anonymous):

hello does anyone want to help

rishavraj (rishavraj):

option B .....is it \[\frac{ 1 }{ x^{-1} }~~~or~~~ \frac{ 1 }{ x - 1 }\]

rishavraj (rishavraj):

see if Option B is \[\frac{ 1 }{ x^{-1} }\] then u must be knowing tht \[\frac{ 1 }{ x^{-1} } = x^1 = x\] and see option D i.e \[x^{\frac{ 1 }{ 3 }} \times x^{\frac{ 1 }{ 3 }} \times x^{\frac{ 1 }{ 3 }} ~~=~~ x^{\frac{ 1 }{ 3 } + \frac{ 1 }{ 3 } + \frac{ 1 }{ 3 }} = x^\frac{ 3 }{ 3 } = x\]